# Accelerated ADMM: Automated Parameter Tuning and Improved Linear Convergence

### v1

@@ 5 Case Study
We test the A-ADMM schemes designed based on the theoretical analyses in the previous section on the $\ell_{1}$-regularized least-squares problem, also known as LASSO Regression, and compare them with available algorithms (Boyd et al., 2011). We use the benchmark problem
@@
where $F\in\mathbb{R}^{250\times 100}$ is chosen as a tall and full column rank matrix with entries that are sampled i.i.d. from an isotropic Gaussian distribution $\mathcal{N}(0,1)$, with normalized columns. The vector $b$ is generated as $b=Fw^{0}+\varepsilon$, where $w^{0}\in\mathbb{R}^{100}$ is sparse, containing 50 non-zeros drawn from $\mathcal{N}(0,1)$ and $\varepsilon\sim\mathcal{N}(0,10^{-3}I_{250})$. The regularization parameter $\tau>0$, is set to $\tau=0.01$. Problem (23) fits the general problem (1), where the first term is $m$-strongly convex and $L$-smooth with $m=\underline{\sigma}^{2}(F^{\top}F)$ and $L=\bar{\sigma}^{2}(F^{\top}F)$.
@@
In particular, we study the convergence speed of A-ADMM and OR-A-ADMM, both with TM and GS parameters, against the widely used LASSO benchmark solver FISTA (Beck and Teboulle, 2009) as well as ADMM and OR-ADMM.
@@
Figure 5: Empirical performance of different A-ADMM schemes and other benchmark algorithms applied to LASSO regression.
As shown in Figure 5, the A-ADMM methods with proposed tunings consistently achieve faster convergence compared to both the vanilla ADMM and existing accelerated algorithms from the literature. Within our family of methods, the OR-variant provides a clear advantage as it always outperforms the nominal counterparts. The grid search parameter set yields the best result. We also observe that the TM version with and without $\lambda$-damping show almost identical empirical performance. Notably, the ordering of algorithms evaluated on the empirical performance and evaluated on the worst-case convergence rates stays the same, suggesting that the worst-case convergence rate is a suitable metric for algorithm comparison.
@@ 6 Conclusions
This paper used the IQC-based algorithm analysis framework to show linear convergence rates of different accelerated ADMM variants. We proposed different parameter tunings and show that all variants outperform vanilla ADMM. The fastest version both in terms of certified convergence rate and empirical performance was found using a grid search over the parameter space, giving the fastest version of A-ADMM in the strongly-convex-smooth setting to date. We moreover showed the close connection between A-ADMM parameters and accelerated gradient methods from unconstrained optimization. Future works could focus on theoretic derivations of the optimal A-ADMM parameters with tools from systems theory, and the development of a systematic and robust A-ADMM synthesis procedure.

### v2

@@ 5 Case Study
We test the A-ADMM schemes designed based on the theoretical analyses in the previous section on the $\ell_{1}$-regularized least-squares problem, also known as LASSO Regression, and compare them with available algorithms (boydAdmm). We use the benchmark problem
@@
where $F\in\mathbb{R}^{500\times 100}$ is chosen as a tall and full column rank matrix with entries that are sampled i.i.d. from an isotropic Gaussian distribution $\mathcal{N}(0,1)$, with normalized columns. The vector $b$ is generated as $b=Fw^{0}+\varepsilon$, where $w^{0}\in\mathbb{R}^{100}$ is sparse, containing 50 non-zeros drawn from $\mathcal{N}(0,1)$ and $\varepsilon\sim\mathcal{N}(0,10^{-3}I_{500})$. The regularization parameter $\tau>0$, is set to $\tau=0.01$. Problem (23) fits the general problem (1), where the first term is $m$-strongly convex and $L$-smooth with $m=\underline{\sigma}^{2}(F^{\top}F)$ and $L=\bar{\sigma}^{2}(F^{\top}F)$.
@@
In particular, we study the convergence speed of A-ADMM and OR-A-ADMM, both with TM and GS parameters, against the widely used LASSO benchmark solver FISTA (FISTA) as well as ADMM and OR-ADMM. The iterate error computed on one random instance of (23) is shown in Fig. 5, while Table 3 summarizes the average number of iterations required to reach $\Delta_{k}<10^{-5}$ over 200 random instances.
@@
Figure 5: Empirical performance of different A-ADMM schemes and other benchmark algorithms applied to a random instance of LASSO regression.
Table 3: Average iterations to reach $\Delta_{k}<10^{-6}$ over 200 random LASSO instances.
Method Average iterations
ADMM 46.19
A-ADMM (TM) 46.24
FISTA 39.3
OR-ADMM 37.48
A-ADMM (GS) 36.66
OR-A-ADMM (GS) 34.98
Both Fig. 5 and Table 3 show that the performance of the A-ADMM methods differ profoundly based on their parameter selection. While a TM tuning is empirically showing no advantage over an optimally tuned vanilla ADMM, we see that a GS tuning gives a substantial advantage. In addition, we observe that the OR-ADMM with GS tuning consistently outperforms all other methods. Notably, the ordering of algorithms evaluated on the empirical performance and evaluated on the worst-case convergence rates stays the same, suggesting that the worst-case convergence rate is a suitable metric for algorithm comparison.
@@ 6 Conclusions
This paper used the IQC-based algorithm analysis framework to show linear convergence rates of different accelerated ADMM variants. We proposed different parameter tunings, and show that the additional momentum parameters allows to outperform vanilla ADMM. The fastest version both in terms of certified convergence rate and empirical performance was found using a grid search over the parameter space, giving the fastest version of A-ADMM in the strongly-convex-smooth setting to date. We moreover showed the close connection between A-ADMM parameters and accelerated gradient methods from unconstrained optimization. Future works could focus on theoretic derivations of the optimal A-ADMM parameters, and the development of a systematic and robust A-ADMM synthesis procedure.